### v1

# Ranking matters: Does the new format select the best teams for the knockout phase in the UEFA Champions League?

Abstract
Starting in the 2024/25 season, the Union of European Football Associations (UEFA) has fundamentally changed the format of its club competitions: the group stage has been replaced by a league phase played by 36 teams in an incomplete round robin format. This makes ranking the teams based on their results challenging because teams play against different sets of opponents, whose strengths vary. In this research note, we apply several well-known ranking methods for incomplete round robin tournaments to the 2024/25 UEFA Champions League league phase. Our results show that it is doubtful whether the currently used point-based system provides the best ranking of the teams.
Keywords: fairness; incomplete round robin tournament; ranking; strength of schedule; UEFA Champions League
MSC class: 15A06, 91B14
JEL classification number: C44, Z20
1 Introduction
The 2024/25 UEFA Champions League has introduced a new format, replacing the traditional group stage with an incomplete round robin tournament where each of the 36 teams plays against eight different opponents. The official ranking is a lexicographical order based on the number of points, followed by goal difference, and the number of goals scored as tiebreakers. While using four strength-based pots is intended to create opponent sets of comparable strength, the difficulty of a team’s set of opponents – known as strength of schedule (SoS), see, e.g. Fearnhead and Taylor, (2010) – can still vary considerably between different teams.
Table 1: Strength of schedules in the 2024/25 UEFA Champions League league phase
Team Pot 1 Pot 2 Pot 3 Pot 4 Sum SoS
Liverpool 15 3 16 15 16 14 6 3 88 85
Barcelona 15 15 15 13 0 6 13 13 90 86
Arsenal 13 19 7 15 11 11 13 3 92 88
Inter Milan 3 11 19 16 6 0 13 4 72 68
Atleti 3 13 16 13 16 3 0 4 68 62
Leverkusen 19 21 15 18 3 13 4 13 106 99
Lille 15 21 12 18 13 11 6 6 102 95
Aston Villa 15 3 12 11 12 0 6 13 72 65
Atalanta 15 19 19 7 12 0 6 10 88 82
B. Dortmund 19 15 7 11 12 11 6 6 87 78
Real Madrid 15 21 15 15 3 16 10 13 108 99
Bayern München 13 19 13 7 11 13 0 16 92 83
Milan 21 15 11 16 6 11 3 0 83 74
PSV 21 13 7 12 11 6 3 13 86 78
Paris 11 15 18 19 14 3 3 10 93 83
Benfica 19 15 18 12 13 6 6 13 102 92
Monaco 19 19 13 19 6 11 16 6 109 99
Brest 15 19 16 7 14 3 6 4 84 74
Feyenoord 15 11 16 13 3 16 4 3 81 71
Juventus 11 3 13 11 14 16 10 16 94 85
Celtic 3 15 11 15 0 11 0 16 71 62
Man City 19 13 11 12 13 11 4 0 83 72
Sporting CP 11 3 19 11 16 14 6 6 86 75
Club Brugge 15 11 12 15 11 12 16 6 98 87
GNK Dinamo 15 15 15 19 12 3 13 0 92 81
Stuttgart 13 15 15 12 0 6 4 0 65 52
Shakhtar 15 15 15 19 0 14 13 6 97 81
Bologna 15 21 7 13 16 11 13 16 112 97
Crvena Zvezda 19 19 13 15 14 0 10 13 103 85
Sturm Graz 3 15 11 15 11 16 3 13 87 69
Sparta Praha 19 11 18 16 3 13 13 10 103 84
Leipzig 21 19 12 18 11 12 16 6 115 94
Girona 21 13 19 15 13 14 0 6 101 80
Salzburg 13 15 18 16 11 13 13 4 103 82
S. Bratislava 11 15 15 18 11 12 10 3 95 71
Young Boys 19 19 15 7 6 12 16 10 104 80
  • The column Sum shows the number of points scored by the eight opponents.
  • The column SoS is the sum of opponents’ points minus the sum of points lost by the team against them.
This is evident from Table 1, which presents the number of points obtained by the opponents of each team (two from each pot). We calculate the strength of schedule of a team as the sum of its opponents’ points, minus the points the team itself lost against them. Notably, Man City and Sporting CP, both of whom survived the league stage, faced significantly weaker opponents than GNK Dinamo that obtained the same number of points but was eliminated. Furthermore, GNK Dinamo was ranked below these teams based on goal difference, which was highly negative primarily due to their heavy defeat (2-9) against Bayern München in the first round. However, teams that face weaker opponents have an advantage regarding goal difference since securing high-margin victories is easier against weaker opponents, which again raises questions of fairness. Analogously, in the race for the Round of 16, placing Aston Villa in the Top 8 appears to have been strongly influenced by its relatively easy schedule. This stands in contrast to teams like Atalanta, Borussia Dortmund, Real Madrid, Bayern München, and Milan, all of them scoring just one point less.
The current note aims to explore alternative ranking methods that provide a different perspective on team performances by taking the strength of the schedule into account. By applying these approaches to the final league-phase standing in the 2024/25 UEFA Champions League, we illustrate how team rankings vary considerably depending on the chosen methodology. As the final ranking determines which teams can qualify for the next phase, these results show the non-negligible effect of the ranking method on the fairness of the competition.
2 Alternative ranking methods
As an alternative to the official point-based ranking, we examine three ranking methods suggested for incomplete round-robin tournaments, where teams face varying strengths of schedule: Keener’s direct ranking method (Section 2.1), the generalized row sum (Section 2.2), and Colley’s ranking (Section 2.3). Each of these ranking methods is derived by solving a system of linear equations that gives the strength $r_{i}$ of each team $i$. Note that other ranking methods exist, see, e.g. Devlin and Treloar, (2018), Vaziri et al., (2018), and Dabadghao and Vaziri, (2022).
2.1 Keener’s direct ranking method
Keener’s ranking method assumes that the ranking is proportional to both the performance of teams against each of its opponents, as well as to the strength of its opponents. This ranking is obtained by solving the following linear system $Ar=\lambda r$:
$\begin{pmatrix}\frac{a_{11}}{t_{1}}&\frac{a_{12}}{t_{1}}&\dots&\frac{a_{1n}}{t_{1}}\\ \frac{a_{21}}{t_{2}}&\frac{a_{22}}{t_{2}}&\dots&\frac{a_{2n}}{t_{2}}\\ \dots&\dots&\dots&\dots\\ \frac{a_{n1}}{t_{n}}&\frac{a_{n2}}{t_{n}}&\dots&\frac{a_{nn}}{t_{n}}\\ \end{pmatrix}\begin{pmatrix}r_{1}\\ r_{2}\\ \dots\\ r_{n}\end{pmatrix}=\lambda\begin{pmatrix}r_{1}\\ r_{2}\\ \dots\\ r_{n}\end{pmatrix}$ (1)
where $t_{i}$ is the number of games played by team $i$ (in our case $t_{i}=8$ for all $i$), and $a_{ij}$ is the result of match between teams $i$ and $j$ for team $i$. In particular, $a_{ij}=3$ if $i$ wins, $a_{ij}=1$ if the result is a draw, and $a_{ij}=0$ if $i$ loses. The vector $(r_{1},r_{2},\dots,r_{n})^{T}$ represents the strength ratings of the teams, which are used to determine the ranking. Note that this is an eigenvector of $A$ and $\lambda$ is the associated eigenvalue. When applied to a complete round robin tournament where each team plays the same number of matches against all other teams, the direct ranking method is not necessarily the same as the point-based ranking. More information can be found in Keener, (1993) and Lambers and Spieksma, (2020).
2.2 Generalized Row Sum
The generalized row sum (GRS) is a parametric family of ranking methods, which adjusts the average number of points per game by taking into account the performance of the opponents. Specifically, a ranking is obtained by solving the following linear system:
$\left[\begin{pmatrix}1&0&\dots&0\\ 0&1&\dots&0\\ \dots&\dots&\dots&\dots\\ 0&0&\dots&1\end{pmatrix}+\varepsilon\begin{pmatrix}t_{1}&-n_{12}&\dots&-n_{1n}\\ -n_{21}&t_{2}&\dots&-n_{2n}\\ \dots&\dots&\dots&\dots\\ -n_{n1}&-n_{n2}&\dots&t_{n}\end{pmatrix}\right]\begin{pmatrix}r_{1}\\ r_{2}\\ \dots\\ r_{n}\end{pmatrix}=\begin{pmatrix}s_{1}\\ s_{2}\\ \dots\\ s_{n}\end{pmatrix}$ (2)
where $s_{i}$ is the normalized number of points of team $i$, $\varepsilon\geq 0$ is a parameter, and $n_{ij}=n_{ji}$ is the number of matches between teams $i$ and $j$ (in our case, this is either 0 or 1). The scores are normalized as follows. First, the average score per match is calculated for each team by dividing its number of points by the number of matches played. Then, these values are normalized by subtracting the average quotient from the quotient of each team to obtain $s_{i}$.
The first term corresponds with the number of points obtained by the teams, while the second term reflects their strength of schedule. If $\varepsilon=0$, this method is equivalent to the points-based ranking without any tie-breaking criterion. Strength of schedule can be used as a tie-breaking rule if the value of parameter $\varepsilon$ is small. However, higher values of parameter $\varepsilon$ may swap teams that have different numbers of points. More details about this method can be found in Chebotarev, (1994), González-Díaz et al., (2014), Csató, (2017) and Csató, (2021).
A crucial property of the generalized row sum method is that it implies the same ranking for any value of $\varepsilon$ as the number of points if each team plays the same number of games against any other team, namely, if the tournament is round robin. Leiva Bertrán, (2025) has recently proposed a new family of scoring methods for ranking teams in incomplete tournaments, which satisfies a number of reasonable theoretical properties (e.g. win dominance, win/loss fairness). This method coincides with the generalized row sum if all teams play the same number of matches (Leiva Bertrán,, 2025, Proposition 1) as in the UEFA Champions League league phase.
2.3 Colley
Colley’s ranking method aims to rank the teams based on their winning percentage and the strength of the schedule. This ranking is obtained by solving the following linear system of equations:
$\begin{pmatrix}2+t_{1}&-n_{12}&\dots&-n_{1n}\\ -n_{21}&2+t_{2}&\dots&-n_{2n}\\ \dots&\dots&\dots&\dots\\ -n_{n1}&-n_{n2}&\dots&2+t_{n}\end{pmatrix}\begin{pmatrix}r_{1}\\ r_{2}\\ \dots\\ r_{n}\end{pmatrix}=\begin{pmatrix}b_{1}\\ b_{2}\\ \dots\\ b_{n}\end{pmatrix},$ (3)
where $b_{i}=1+\left(w_{i}-\ell_{i}\right)/2$ with $w_{i}$ being the number of wins plus half the number of draws, and $\ell_{i}$ being the number of losses plus half the number of draws. More information can be found in Colley, (2002), Chartier et al., (2011) and Vaziri et al., (2018).
3 Results
Table 2: Rankings with three alternative methods in the 2024/25 UEFA Champions League league phase
Method Points Official ranking Keener GRS ($\varepsilon=0.01$) GRS ($\varepsilon=\infty$) Colley
Liverpool 21 1 1 1 1 1
Barcelona 19 2 2 3 2 4
Arsenal 19 3 3 2 3 2
Inter Milan 19 4 5 4 6 3
Atleti 18 5 10 5 8 5
Leverkusen 16 6 6 6 5 8
Lille 16 7 4 7 4 9
Aston Villa 16 8 13 8 15 7
Atalanta 15 9 18 11 14 6
B. Dortmund 15 10 14 12 12 18
Real Madrid 15 11 7 9 7 15
Bayern München 15 12 11 10 10 16
Milan 15 13 17 13 13 11
PSV 14 14 12 14 16 13
Paris 13 15 22 17 18 20
Benfica 13 16 8 16 11 12
Monaco 13 17 9 15 9 10
Brest 13 18 21 18 17 17
Feyenoord 13 19 16 19 19 24
Juventus 12 20 15 20 20 14
Celtic 12 21 25 21 25 25
Man City 11 22 23 25 24 23
Sporting CP 11 23 19 24 23 21
Club Brugge 11 24 20 22 22 19
GNK Dinamo 11 25 24 23 21 22
Stuttgart 10 26 27 26 29 26
Shakhtar 7 27 28 27 27 29
Bologna 6 28 26 28 26 27
Crvena Zvezda 6 29 30 29 28 28
Sturm Graz 6 30 32 30 30 33
Sparta Praha 4 31 33 31 32 31
Leipzig 3 32 31 32 31 30
Girona 3 33 34 34 34 34
Salzburg 3 34 29 33 33 32
S. Bratislava 0 35 35 36 36 36
Young Boys 0 36 36 35 35 35
  • The second column Ranking shows the position in the official league table.
  • GRS stands for Generalized row sum.
The rankings produced by the alternative ranking methods in Section 2 are shown in Table 2. For the generalized row sum, we provide the two extremal results: when the strength of schedule is used only for tie-breaking ($\varepsilon$=0.01) and when the performance of the opponents has the highest possible effect ($\varepsilon=\infty$). The rankings for other values of the parameter $\varepsilon$ can be found in the Appendix, in Table A.1.
Clearly, different methods produce different rankings. With respect to qualification, the clubs are essentially allocated into three sets: ranked in Top 8 (direct qualification for the Round of 16), ranked between 9th and 24th (qualification for the knockout phase play-offs), and ranked between 25th and 36th (elimination). In view of these thresholds, the results can be summarized as follows:
  • Aston Villa falls out of the Top 8 when the strength of schedule has a higher impact. In particular, it only ranks 13th in Keener’s direct ranking method and even only 15th in the GRS with $\varepsilon=\infty$. On the other hand, Real Madrid is ranked 7th according to Keener’s method and GRS ($\varepsilon=\infty$). Hence, it can be argued that Real Madrid should have directly qualified for the Round of 16 at the expense of Aston Villa, even though the latter club has scored one point more.
  • Monaco is clearly the best team among the five with 13 points if the schedule of strength is taken into account, and it can even be ranked higher than some teams with 15 points. Similarly, Benfica could also be ranked higher; it is placed in the Top 8 by Keener’s method.
  • All alternative rankings except GRS if the strength of schedule plays a minor role ($\varepsilon<0.5$) exclude Celtic from the Top 24. However, the Croatian champion GNK Dinamo is eliminated only according to the official ranking rule; if its strength of schedule is taken into account, it is never ranked below the 23rd position.
Nonetheless, some teams are consistently ranked in the same position:
  • The official ranking, GRS and Keener’s method agree undeniably on the top three teams, which are Liverpool, Barcelona, and Arsenal in this order.
  • In the GRS, seven teams have a guaranteed place in the Top 8 independently of the value of parameter $\varepsilon$.
  • All ranking methods agree on the elimination of nine teams (Stuttgart, Shakhtar, Crvena Zvezda, Sturm Graz, Sparta Praha, Girona, Salzburg, S. Bratislava and Young Boys).
In some cases, the ranking methods produce remarkably different results. For example, Atalanta is officially ranked 9th, while it is ranked 18th according to Keener’s direct method. The result of the Colley ranking method is quite similar to the official ranking. As for the GRS ranking, playing only eight rounds in a league with 36 teams seems to be insufficient to obtain a unique ranking. In contrast, in the 2019/20 season of the double round robin European football leagues, where 18–20 teams played 25–29 rounds before the COVID-19 pandemic, the ranking with the generalized row sum method was essentially insensitive to the value of parameter $\varepsilon$ (Csató,, 2021). Indeed, playing a higher number of rounds increases the variance of the number of points and decreases the variance in the strength of opponents, reducing the probability that a team with fewer points scored against stronger opponents is ranked above a team with more points scored against weaker opponents.
4 Discussion
Ranking teams in an incomplete round robin tournament is challenging since it is difficult to choose an uncontroversial balance between the number of points scored and the strength of schedule. Our results reveal that this issue is also relevant for the 2024/25 UEFA Champions League league phase: both the set of teams directly qualified for the Round of 16 and the set of teams eliminated would have changed when using reasonable alternative ranking methods. Consequently, the current mechanism using four pots based on strengths according to the UEFA club coefficients may not sufficiently reduce the impact of different strengths of opponent sets.
Possible remedies include (1) using a ranking that reflects the strengths of the teams better than UEFA club coefficients to assign the teams to pots (e.g. a ranking based on Elo ratings as demonstrated by Csató, (2024)), and (2) using the strength of schedule as the primary tie-breaking criterion instead of goal difference (which would boil down to the GRS ranking with, for example, $\epsilon=0.01$). Both recommendations seem to be realistic since the Elo approach has been adopted for the FIFA World Ranking in 2018 (FIFA,, 2018) and the Buchholz method, based on the arithmetic mean of the opponents points’, is the most common tie-breaking rule in Swiss-system tournaments (Freixas,, 2022).
Acknowledgements
The research was supported by the National Research, Development and Innovation Office under Grants FK 145838, and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.
Appendix
Table A.1: Rankings with the Generalized Row Sum method in the 2024/25 UEFA Champions League league phase
Parameter ($\varepsilon$) Ranking 0 0.01 0.1 0.25 0.5 0.75 1 2 5 10 100 $\infty$ Liverpool 1 1 1 1 1 1 1 1 1 1 1 1 1 Barcelona 2 2 3 3 2 2 2 2 2 2 2 2 2 Arsenal 3 2 2 2 3 3 3 3 3 3 3 3 3 Inter Milan 4 2 4 4 4 4 4 4 5 6 6 6 6 Atleti 5 5 5 5 7 8 8 8 8 8 8 8 8 Leverkusen 6 6 6 6 5 5 6 6 6 5 5 5 5 Lille 7 6 7 7 6 6 5 5 4 4 4 4 4 Aston Villa 8 6 8 9 11 13 15 15 15 15 15 15 15 Atalanta 9 9 11 12 13 14 14 14 14 14 14 14 14 B. Dortmund 10 9 12 11 10 11 11 11 11 12 12 12 12 Real Madrid 11 9 9 8 8 7 7 7 7 7 7 7 7 Bayern München 12 9 10 10 9 9 10 10 10 10 10 10 10 Milan 13 9 13 13 12 12 12 12 13 13 13 13 13 PSV 14 14 14 15 16 16 16 16 16 16 16 16 16 Paris 15 15 17 17 17 17 17 18 18 18 18 18 18 Benfica 16 15 16 16 15 15 13 13 12 11 11 11 11 Monaco 17 15 15 14 14 10 9 9 9 9 9 9 9 Brest 18 15 18 18 18 18 18 17 17 17 17 17 17 Feyenoord 19 15 19 19 19 19 19 19 19 19 19 19 19 Juventus 20 20 20 20 20 20 20 20 20 20 20 20 20 Celtic 21 20 21 23 24 25 25 25 25 25 25 25 25 Man City 22 22 25 25 25 24 24 24 24 24 24 24 24 Sporting CP 23 22 24 24 23 23 23 23 23 23 23 23 23 Club Brugge 24 22 22 21 21 21 22 22 22 22 22 22 22 GNK Dinamo 25 22 23 22 22 22 21 21 21 21 21 21 21 Stuttgart 26 26 26 26 26 28 28 28 29 29 29 29 29 Shakhtar 27 27 27 27 28 27 27 27 27 27 27 27 27 Bologna 28 28 28 28 27 26 26 26 26 26 26 26 26 Crvena Zvezda 29 28 29 29 29 29 29 29 28 28 28 28 28 Sturm Graz 30 28 30 30 30 30 30 30 30 30 30 30 30 Sparta Praha 31 31 31 31 31 32 32 32 32 32 32 32 32 Leipzig 32 32 32 32 32 31 31 31 31 31 31 31 31 Girona 33 32 34 34 34 34 34 34 34 34 34 34 34 Salzburg 34 32 33 33 33 33 33 33 33 33 33 33 33 S. Bratislava 35 35 36 36 36 36 36 36 36 36 36 36 36 Young Boys 36 35 35 35 35 35 35 35 35 35 35 35 35 Teams are ranked according to the official league phase table. The second column Ranking shows the position in the official league table.